# Machine Learning Assisted Adjustment Boosts Inferential Efficiency of Randomized Controlled Trials

### v1

@@ 4 Diet Intervention Data Example
In this example, the investigators studied whether a brief diet intervention can reduce the symptoms of depression in young adults. There were 76 particpants randomly allocated to a brief 3-week diet intervention (Diet Group) or a habitual diet control group (Control Group), with $n=38$ per group. The primary and secondary outcome measures were assessed at both baseline and after the intervention, which included symptoms of depression (Centre for Epidemiological Studies Depression Scale, CESD-R; and Depression Anxiety and Stress Scale-21 depression subscale, DASS-21-D), current mood (Profile of Mood States, POMS), self-efficacy (New General Self-Efficacy Scale, GSES) and memory (Hopkins Verbal Learning Test). There are a total of 23 baseline variables can be used as covariates. The GSES score were log-transformed as specified in the article. The ANCOVA was used for the analysis of both primary and secondary outcomes with the baseline scores adjusted as covariates. The results showed that the Diet group had significantly lower self-reported depression symptoms than the Control Group on the CESD-R and DASS-21 depression subscale. However, none of the POMS scores showed significant difference. Here we used our method (Wilcoxon-RF) to re-analyzed the difference of POMS fatigue score and two-sided tests were used. The results in Table 1 show that only the Wilcoxon-RF suggest there is a significant difference in the post-intervention fatigue scores between two groups. This demonstrates the higher statistical power of the proposed method. Based on the residuals, the final estimated treatment effect is -1.57, with 95% confidence interval of (-3.01, -0.05). The variable importance from RF suggests baseline depression is the most important predictor. The partial dependence plot (Figure 6) suggests the post-intervention fatigue does not change with the baseline depression within the range of $0\sim 5$, but starts to show a linear trend in the range of $6\sim 12$, and then plateaus. Such sigmoid shape may not be well described by a linear model, but can be better captured by a nonlinear model like RF, which explains why our method achieves better efficiency.
@@ 5 Discussion
In this work, we presented a novel inferential procedure for RCTs assisted by an machine learning adjustment. Through a set of simulation experiments and a real world example, we showed the proposed method remarkably boosts the statistical power and can reduce 1/3 sample size requirement under certain scenarios. Such improvement is grounded in the capability of machine learning in capturing nonlinear relationships and higher order interactions, which cannot be fully exploited by linear models. We also demonstrated that the proposed method is generally safe to use in practice because the type I error is robustly controlled. Meanwhile, under the scenarios tested, it never shows worse performance than the straight two-sample tests without adjustments. In contrary, the methods using adjustment by linear regression can sometimes harm the efficiency, which can be attributed to its instability when $p$ is large relatively to $N$. Shrinkage methods such as Ridge [3], Lasso [4] or Elastic net [6] may be used, but they all require careful tuning of hyperparameters, and they do not address the difficulty brought by non-linearities and interactions.
@@
The RF is often considered an “off-the-shelf” supervised learning method due to its simplicity of usage. Typically with the recommended settings one can obtain a model with good performance [8]. This makes it a good choice for RCTs which typically requires detailed analysis plans to be pre-specified in the study protocols. The usage of OOB observations further makes explicit training, validation set splitting unnecessary. Therefore, in an ideal scenario, the inference procedure can be completed in a single run of RF model training. Another motivation for using RF is that it usually shows better or comparable prediction performance than neural networks in the range of sample sizes typically used in a randomized study. Deep learning models typically require larger sample size to gain advantage, and it usually requires extensive tuning [11], which can complicate the inference procedure and is hard to plan beforehand. However, no model can be optimal under all scenarios. Although RF is used as a proof of concept in our study, other modeling methods can be superior to RF in many settings. Therefore, investigation of the proposed approach to incorporate other machine learning techniques is important.
@@
Another practical challenge is how to translate the gain in efficiency into reduction of sample size. This requires the knowledge of the variation that can be explained by the machine learning model at the planning stage. In medical research, a phase II trial is conducted to preliminarily assess the efficacy before conducting a large scale phase III RCT. The data collected from the phase II trial provides important parameters that can guide the design of the phase III trial. One can simply use the two-sample $t$-test formula as an approximation for the exact test, and estimate the proportion of variance that will be reduced by including the covariates based on assumptions or data available, e.g. from a phase II trial. Approximately, if variance explained by the model is $\gamma$, then one would expect a reduction of $N\gamma$ in sample size requirement, where $N$ is the sample size needed using a unadjusted two-sample test. The parameters may also be used to run simulations to estiamte the sample sizes required for the machine learning adjusted tests. Another option is an interim analysis on the efficiency gain and adjust the sample size accordingly. Such strategy may require further methodological investigation.

### v2

@@ 4 Diet Intervention Data Example
In this example, the investigators studied whether a brief diet intervention can reduce the symptoms of depression in young adults. There were 76 participants randomly allocated to a brief 3-week diet intervention (Diet Group) or a habitual diet control group (Control Group), with $n=38$ per group. The primary and secondary outcome measures were assessed at both baseline and after the intervention, which included symptoms of depression (Centre for Epidemiological Studies Depression Scale, CESD-R; and Depression Anxiety and Stress Scale-21 depression subscale, DASS-21-D), current mood (Profile of Mood States, POMS), self-efficacy (New General Self-Efficacy Scale, GSES) and memory (Hopkins Verbal Learning Test). There are a total of 23 baseline variables that can be used as covariates. The GSES score were log-transformed as specified in the article. The ANCOVA was used for the analysis of both primary and secondary outcomes with the baseline scores adjusted as covariates. The results showed that the Diet group had significantly lower self-reported depression symptoms than the Control Group on the CESD-R and DASS-21 depression subscale. However, none of the POMS scores showed significant difference. Here we used the Wilcoxon-RF method to re-analyze the difference in POMS fatigue scores between two groups. Two-sided tests were used at $\alpha=0.05$. The results in Table 1 show that only the Wilcoxon-RF shows a significant difference in the post-intervention fatigue scores between two groups ($p=0.0451$). This demonstrates the higher statistical power of the proposed method. Based on the residuals, the final estimated treatment effect is -1.57 (Diet vs. Control), with 95% confidence interval of (-3.01, -0.05). The variable importance from RF model suggests baseline depression is the most important predictor. The partial dependence plot (Figure 4) suggests the post-intervention fatigue does not change with the baseline depression within the range of $0\sim 5$, but starts to show a linear trend in the range of $6\sim 12$, and then plateaus. The sigmoid shape of what may not be well described by a linear model, but can be better captured by a nonlinear model like RF, which explains why our method achieves better efficiency.
@@ 5 Discussion
In this work, we presented a novel inferential procedure for RCTs assisted by a machine learning adjustment. Through a set of simulation experiments and a real-world example, we showed the proposed method remarkably boosts the statistical power and can reduce 1/3 sample size requirement as compared with other approaches. Such improvement is grounded in the capability of machine learning in capturing nonlinear relationships and higher order interactions, which cannot be fully exploited by linear models. We also demonstrated that the proposed method has proper type I error control. Meanwhile, as long as the outcome has a moderate level of dependency with the baseline covariates, the Wilcoxon-RF method never performs worse than the standard two-sample tests. Hence, the risk of using the proposed method is minimal, while the potential gain is substantial. In contrary, the methods using adjustment by linear regression can sometimes harm the efficiency, which can be attributed to its instability when $p$ is large relatively to $N$. Shrinkage methods such as Ridge [6], Lasso [7] or Elastic net [9] may be used, but they all require careful tuning of hyperparameters, and they do not address the difficulty brought by non-linearities and interactions.
@@
The RF model is often considered an “off-the-shelf” supervised learning method due to its simplicity of usage. Typically, with the recommended settings one can obtain a model with good performance [11]. The usage of OOB observations further makes explicit training, validation set splitting unnecessary. Therefore, in an ideal scenario, the inference procedure can be completed in a single run of RF model training. This feature makes it particularly advantageous because the FDA has clear expectations that the analyses of primary and secondary endpoints in clinical trials need to be pre-specified. The post-hoc analyses are prone to data-driven manipulations which can lead to biased interpretation of the results. Therefore, any modeling method that is difficult to preplan and requires extensive tuning of hyperparameters can negatively affect the weight of evidence from an RCT for the regulatory decision-making. Another motivation for using RF is that it usually shows better or comparable prediction performance than neural networks in the range of sample sizes typically used in a randomized study. Deep learning models usually require larger sample size to gain advantage [12]. However, no model can be optimal under all scenarios. Although RF is used as a proof of concept in our study, other modeling methods can be superior to RF in many settings. Therefore, investigation of the proposed approach to incorporate other machine learning techniques is important.
@@
Another practical challenge is how to translate the gain in efficiency into a reduction of the sample size. This requires the knowledge of the variation that can be explained by the machine learning model at the planning stage. In medical research, a phase II trial is conducted to preliminarily assess the efficacy before conducting a large scale phase III RCT. The data collected from the phase II trial provides important parameters that can guide the design of the phase III trial. One can simply use the two-sample $t$-test formula as an approximation for the exact test, and estimate the proportion of variance that will be reduced by including the covariates based on assumptions or data available, e.g. from a phase II trial. Approximately, if variance explained by the model is $\gamma$, then one would expect a reduction of $N\gamma$ in sample size requirement, where $N$ is the sample size needed using an unadjusted two-sample test. The parameters may also be used to run simulations to estimate the sample sizes required for the machine learning adjusted tests. Another option is an interim analysis on the efficiency gain and adjust the sample size accordingly. Such strategy may require further methodological investigation.